OpenStudy (nuccioreggie):

Convert 3,200,000 to scientific notation. 3.2 ⋅ 106 3.2 ⋅ 105 3.2 ⋅ 10−6 3.2 ⋅ 10−5

OpenStudy (nuccioreggie):

@LegendarySadist

OpenStudy (anonymous):

\[\large 3,200,200.0\] Move the decimal to the left until you get it between the 3 and 2 :)

OpenStudy (nuccioreggie):

@LegendarySadist

OpenStudy (anonymous):

And that should be a 0, not a second 2.

OpenStudy (nuccioreggie):

Yea i think its C

OpenStudy (anonymous):

nooo

OpenStudy (nuccioreggie):

Yea i think its C

OpenStudy (nuccioreggie):

What how

OpenStudy (nuccioreggie):

What how

OpenStudy (nuccioreggie):

What how

OpenStudy (anonymous):

\[\large 3.2 \times 10^{-6} = \frac{1}{3.2 \times 10^{6}}\]

OpenStudy (anonymous):

Let's work backwards. Say we have 3.200000 How many times would we have to move the decimal to the right to get 3,200,000?

OpenStudy (nuccioreggie):

Oooooohok ur right i did it wrong it was A i shouldv known

OpenStudy (anonymous):

Yep, it would be A :)